OpenStudy (anonymous):

Use the dot product to find the magnitude of u if u = 2i - 3j.

OpenStudy (anonymous):

\[\left| u^2 \right|=\left( vector u \right)^2=u.u=\left( 2i-3j \right).\left( 2i-3j \right)=2\left( 2 \right)+\left( -3 \right)\left( -3 \right)=13\] \[\left| u \right|=\sqrt{13}\]

OpenStudy (anonymous):

in the beginning \[\left| u \right|^2=u^2\]